If triangle ABC is right angled at b the sin c+ cos c is bigger or smaller or equal to one ? And why ?

Grade:9

1 Answers

Darsh patel
6 years ago
First of all you should be aware that, for a triangle , the property is that the sum of any two side should be greater than or equal to the third side than only a triangle can be formed. Now in ∆ABC , you can observe that. SinC =AB/AC. and CosC= BC/AC. Now Sin C +Cos C= (AB+BC)/AC. And as I told you that sum of any two side should be greater than or equal to the third side so (AB+BC)≥AC , So Sin C + Cos C is greater than or equal to 1..........Sorry for the inconvenience in the writing as the software of askIitians is not updated and so plz try to understand what I have written.. If not draw a triangle ABC and then verify my solution. Also plz approve my answer......with regards -Darsh Payel